# A Two-Timescale Primal-Dual Framework for Reinforcement Learning via Online Dual Variable Guidance

### v2

@@ 6 Concluding reflections
We proposed a new LP-based primal-dual RL algorithm for entropy-regularized MDPs, suing a single on-policy trajectory generated by a behavioral policy that may evolve with the dual iterate. Using the dynamical-systems approach to stochastic approximation [Borkar, 2023], we establish almost sure last-iterate convergence of the primal and dual iterates to the optimal regularized value function and policy, respectively, under an asynchronous two-timescale scheme with diminishing stepsizes. To the best of our knowledge, such almost sure guarantees have not previously been established for LP-based methods in the entropy-regularized setting. Under a strengthened ergodicity condition, we additionally derive a finite-time last-iterate guarantee, showing that the expected dual objective gap, and consequently the dual mean-squared error, decays at rate $\widetilde{\mathcal{O}}(k^{-2/3})$. This matches the best-known last-iterate behavior for two-timescale stochastic approximation under Markovian sampling [Zeng et al., 2024], despite replay-buffer bias and asynchronous, visitation-adjusted updates along a single trajectory.
@@
While this work focuses on the tabular setting, extending PGDA-RL to function approximation remains a pressing challenge. One promising route is to combine our primal-dual viewpoint and stochastic approximation analysis with LP-based frameworks under linear function approximation, as studied for offline settings by Gabbianelli et al. [2024], and to develop corresponding finite-time guarantees under Markovian single-trajectory data. Such extensions will likely require reformulating the asynchronous update scheme and the structured experience replay mechanism to accommodate approximation error and stability issues. Finally, it is an interesting open question to integrate acceleration ideas for convexified regularized MDP formulations, which have been shown to speed up convergence under full model knowledge [Li et al., 2024], into the present Markovian, replay-based setting.
Acknowledgments.
The authors were supported by the German Research Foundation (DFG) under the grant “SU 1433/2-1".
@@ Appendix A Synchronous Almost Sure Convergence: Proof Details
This section contains the proofs of Propositions 3.1 and 3.2 that are at the heart of the almost sure convergence proof of Algorithm 1.
Proof of Proposition 3.1
@@
We next address the property (A4). The projected dynamic system (13b) has to be well-posed to make claims about its stability. The Lipschitz continuity of the gradient $\nabla_{\rho}L(\lambda(\cdot),\cdot)$ follows from the linearity of the $\lambda(\rho)$ and the Lipschitz continuity of $\nabla_{\rho}L(V,\cdot)$ on the compact set $H$. To clarify the properties of the projection term in (13b), we present the derivation of the projected dynamic following Dupuis [1987]. Define the projection of vector $v\in\mathbb{R}^{|\mathcal{S}|\cdot|\mathcal{A}|}$ at $\rho\in H$ as
@@
Let $N(\rho)$ be the set-valued function that gives the outward normals at $\rho\in\partial H$, that is, $N(\rho)=\{\zeta\in\mathbb{R}^{|\mathcal{S}|\cdot|\mathcal{A}|}\ |\ \forall\rho^{\prime}\in H:\langle\zeta,\rho-\rho^{\prime}\rangle\geq 0,\|\zeta\|=1\}$. Then, by [Dupuis, 1987, Lemma 4.6], the system (34) equals
@@
The existence and uniqueness of a solution to the projected dynamical system are guaranteed by Dupuis and Nagurney [1993, Theorem 2 (uniqueness), Theorem 3 (existence)]. Note that due to the discontinuity of the dynamics, a solution is defined as an absolutely continuous function that fulfills the ODE save on a set of Lebesgue measure zero. To prove the global asymptotic stability of the projected ODE (13b), we next provide a Lyapunov argument based on (35).
@@
serves as a strict Lyapunov function. The inequality $\phi(\rho(t))\geq 0$ holds since $L(\lambda(\rho),\rho)=\min_{V\in\mathbb{R}^{|\mathcal{S}|}}L(V,\rho)\leq\max_{\rho\in H}\min_{V\in\mathbb{R}^{|\mathcal{S}|}}L(V,\rho)=L(V^{\star},\rho^{\star})$. The equality only holds for $\rho=\rho^{\star}$ due to the unique saddle-point as established in Corollary 2.1. Define $\Gamma(\rho(t)):=\nabla_{\rho}L(\lambda(\rho(t)),\rho(t))$ and with a slight abuse of notation $\zeta(\rho(t)):=\zeta(\rho(t),\Gamma(\rho(t)))$. Then,
@@
Lastly, we address the stability of the iterate sequence, property (A5). The dual variable is constrained to the compact set $H$. We, therefore, only have to ensure that $\sup_{k}\|V_{k}\|<\infty$ almost surely. We apply the scaling limit method outlined in Borkar and Meyn [2000] and Borkar [2023, Theorem 4.1], which states that the global asymptotic stability of the origin for a certain scaling limit ODE is a sufficient condition for the almost sure boundedness of the iterates. The scaling limit for the value iterates for a fixed occupancy $\rho$ is defined as
@@ Appendix B Asynchronous Almost Sure Convergence: Proof Details
This section contains the proofs of the asynchronous algorithm’s properties listed in Proposition 4.1. The proofs are presented in a series of lemmas. For (B4) the proof of the synchronous convergence analysis still applies. The two options regarding the behavioral policy update only need to be addressed separately for property (B5).
Lemma B.1.
@@
Proof.
We first derive the conditional expectation of the buffer-based transition estimator. Fix $(s,a)\in\mathcal{D}_{\mathrm{inc}}(s_{k})$, such that at iteration $k$ there is a draw from the corresponding buffer (otherwise $\widehat{\mathcal{P}}_{k}(\cdot|s,a)\equiv 0$ by convention and the claims are trivial). Recall that $\xi_{k}(s,a)$ is drawn uniformly from the list $\mathcal{D}_{k}(s,a)$ and $\widehat{\mathcal{P}}_{k}(s^{\prime}|s,a)=\mathbf{1}\{\xi_{k}(s,a)=s^{\prime}\}$. If $(s,a)=X_{k}=(s_{k-1},a_{k-1})$, then $\mathcal{D}_{k}(s,a)=\mathcal{D}_{k-1}(s,a)\cup\{s_{k}\}$ and $\nu_{k}(s,a)=\nu_{k-1}(s,a)+1$. Conditionally on $\mathcal{F}_{k}^{-}=\sigma(\mathcal{F}_{k-1},s_{k},a_{k})$, the multiset $\mathcal{D}_{k}(s,a)$ is fixed and the uniform draw satisfies
@@
The bounds on the second moments of the martingale differences, (B1)-b, can be derived similarly to the proof of Proposition 3.1 upon noting that $\|\widehat{\mathcal{P}}_{k}(\cdot|s,a)\|_{2}\leq 1$ and
@@
Moreover, boundedness of $\rho_{k-1}$ and $\|\mathcal{P}_{\mathcal{D}_{k}}(\cdot|s,a)\|_{2}\leq 1$ imply that $\{\mathcal{E}_{k}\}$ is uniformly bounded. This proves (B1)-c.
Lemma B.2.
@@
Proof.
The properties (B2) and (B3) regarding the equilibria of the dynamical systems, hold for the asynchronous projected stochastic ascent-descent scheme because we can relate the asymptotic behavior of any solution to the non-autonomous rescaled ODEs (18) to the limiting ODEs of the synchronous setting (13) following [Borkar, 2023, Section 6.4].
@@
Let $\rho\in H$ be a fixed arbitrary dual variable. Let $\epsilon>0$ and $\Lambda^{(1)}(\cdot)$ be a measurable function with $\Lambda^{(1)}(t)\in\Omega^{\epsilon}_{|\mathcal{S}|}$ for all $t$. By definition, $\Lambda^{(1)}(t)$ is a diagonal matrix for each $t$ with bounded diagonal entries $w_{i}(t)\in[\epsilon,1]$. The measurability and boundedness of $t\mapsto\Lambda^{(1)}(t)$ together with the Lipschitz continuity of $V\mapsto\nabla_{V}L(V,\rho)$ imply that the vector field $\Lambda^{(1)}(t)\nabla_{V}L(V,\rho)$ is measurable in $t$ and locally Lipschitz in $V$. Hence, for every initial condition, the ODE (18a) admits a (unique) Carathéodory solution; see, e.g., standard results in differential equations [Filippov, 2013].
@@
Moreover, equality holds if and only if $\nabla_{V}L(V(t),\rho)=0$, that is, $V(t)=\lambda(\rho)$. By $\eta_{V}$-strong convexity of $V\mapsto L(V,\rho)$, the Polyakreplay-buffer biasŁojasiewicz inequality holds: $\|\nabla_{V}L(V,\rho)\|_{2}^{2}\geq 2\eta_{V}\,\Phi_{\rho}(V)$ for all $V$. Therefore,
@@
For (B3) the same arguments hold with the unchanged Lyapunov function used in the proof of condition (A4), see (36), since multiplying the vector field by a diagonal $\Lambda^{(2)}(t)\in\Omega^{\epsilon}_{|\mathcal{X}|}$ preserves negativity of the Lyapunov derivative. ∎
Lemma B.3.
@@
Proof
Let Assumption 1 hold. In Algorithm 2, conditional on the history $\mathcal{F}_{k}$, the next state-action pair $(s_{k+1},a_{k+1})$ is generated by first sampling $s_{k+1}\sim\mathcal{P}(\cdot|s_{k},a_{k})$ and then sampling $a_{k+1}\sim\pi_{k}^{b}(\cdot|s_{k+1})$, where the behavioral policy $\pi_{k}^{b}$ is either fixed or (in the on-policy exploration variant) a deterministic function of $\rho_{k}$ via (3), optionally $\epsilon_{k}$-mixed with the uniform policy. Therefore, for any $(s,a),(s^{\prime},a^{\prime})\in\mathcal{X}$,
@@ Appendix C Asynchronous Convergence Rate: Proof Details
This section contains the proof details of the convergence rate theorem, Theorem 4.2. The first section addresses Lemma 4.1 and Proposition 4.2, followed by the proofs of Propositions 4.3 and 4.4 and finally some additional technical lemmata collected in the third subsection.
C.1 Proofs of Lemma 4.1 and Proposition 4.2
@@
We first address the uniform geometric ergodicity of the Markov chains induced by the dual iterate.
Proof of Lemma 4.1
@@
Let Assumption 3 hold for some deterministic policy $\pi_{d}$. By the equivalence of uniform geometric ergodicity and the multi-step Doeblin condition [Meyn and Tweedie, 2012, Theorem 16.0.2], there exist $M\in\mathbb{N}$, $\epsilon_{0}\in(0,1]$ and a positive probability measure $\upsilon$ such that for the state-transition kernel $\mathcal{P}_{d}$ induced by the deterministic policy $\pi_{d}$ it holds that $\mathcal{P}_{d}^{\,M}(\cdot|s)\geq\epsilon_{0}\,\upsilon(\cdot)$ for all $s\in\mathcal{S}$. Denote $\pi_{\min}:=\tfrac{C^{L}}{|\mathcal{A}|C^{U}}$, then $\pi_{\rho}(a\,|\,s)\geq\pi_{\min}>0$ uniformly over $\rho\in H$, $s\in\mathcal{S}$ and $a\in\mathcal{A}$. Moreover, for any $\rho\in H$, $s,\,s^{\prime}\in\mathcal{S}$, let $\tilde{\mathcal{P}}_{\rho}$ denote the state transition kernel induced by the policy $\pi_{\rho}$. Then
$\tilde{\mathcal{P}}_{\rho}(s^{\prime}|s)\;=\;\sum_{a\in\mathcal{A}}\pi_{\rho}(a\,|\,s)\,\mathcal{P}(s^{\prime}\,|\,s,a)\;\geq\;|\mathcal{A}|\pi_{\min}\,\mathcal{P}(s^{\prime}\,|\,s,\pi_{d}(s))\;=\;|\mathcal{A}|\pi_{\min}\,\mathcal{P}_{d}(s^{\prime}|s).$
@@
By monotonicity under composition, iterating $M$ times yields

### v3

@@ 6 Concluding Reflections
We proposed a new LP-based primal-dual RL algorithm for entropy-regularized MDPs, suing a single on-policy trajectory generated by a behavioral policy that may evolve with the dual iterate. Using the dynamical-systems approach to stochastic approximation (Borkar, 2023), we establish almost sure last-iterate convergence of the primal and dual iterates to the optimal regularized value function and policy, respectively, under an asynchronous two-timescale scheme with diminishing stepsizes. To the best of our knowledge, such almost sure guarantees have not previously been established for LP-based methods in the entropy-regularized setting.
Under a strengthened ergodicity condition, we additionally derive a finite-time last-iterate guarantee, showing that the expected dual objective gap, and consequently the dual mean-squared error, decays at rate $\widetilde{\mathcal{O}}(k^{-2/3})$. This matches the best-known last-iterate behavior for two-timescale stochastic approximation under Markovian sampling (Zeng et al., 2024), despite replay-buffer bias and asynchronous, visitation-adjusted updates along a single trajectory.
@@
While this work focuses on the tabular setting, extending PGDA-RL to function approximation remains a pressing challenge. One promising route is to combine our primal-dual viewpoint and stochastic approximation analysis with LP-based frameworks under linear function approximation, as studied for offline settings by Gabbianelli et al. (2024), and to develop corresponding finite-time guarantees under Markovian single-trajectory data. Such extensions will likely require reformulating the asynchronous update scheme and the structured experience replay mechanism to accommodate approximation error and stability issues. Finally, it is an interesting open question to integrate acceleration ideas for convexified regularized MDP formulations, which have been shown to speed up convergence under full model knowledge (Li et al., 2024), into the present Markovian, replay-based setting.
Acknowledgments and Disclosure of Funding
The authors were supported by the German Research Foundation (DFG) under the grant “SU 1433/2-1”.
@@ Appendix A Synchronous Almost Sure Convergence: Proof Details
This section contains the proofs of Propositions 7 and 8 that are at the heart of the almost sure convergence proof of Algorithm 1.
Proof of Proposition 7
@@
We next address the property (A4). The projected dynamic system (13b) has to be well-posed to make claims about its stability. The Lipschitz continuity of the gradient $\nabla_{\rho}L(\lambda(\cdot),\cdot)$ follows from the linearity of the $\lambda(\rho)$ and the Lipschitz continuity of $\nabla_{\rho}L(V,\cdot)$ on the compact set $H$. To clarify the properties of the projection term in (13b), we present the derivation of the projected dynamic following Dupuis (1987). Define the projection of vector $v\in\mathbb{R}^{|\mathcal{S}|\cdot|\mathcal{A}|}$ at $\rho\in H$ as
@@
Let $N(\rho)$ be the set-valued function that gives the outward normals at $\rho\in\partial H$, that is, $N(\rho)=\{\zeta\in\mathbb{R}^{|\mathcal{S}|\cdot|\mathcal{A}|}\ |\ \forall\rho^{\prime}\in H:\langle\zeta,\rho-\rho^{\prime}\rangle\geq 0,\|\zeta\|=1\}$. Then, by (Dupuis, 1987, Lemma 4.6), the system (34) equals
@@
The existence and uniqueness of a solution to the projected dynamical system are guaranteed by Dupuis and Nagurney (1993, Theorem 2 (uniqueness), Theorem 3 (existence)). Note that due to the discontinuity of the dynamics, a solution is defined as an absolutely continuous function that fulfills the ODE save on a set of Lebesgue measure zero. To prove the global asymptotic stability of the projected ODE (13b), we next provide a Lyapunov argument based on (35).
@@
serves as a strict Lyapunov function. The inequality $\phi(\rho(t))\geq 0$ holds since $L(\lambda(\rho),\rho)=\min_{V\in\mathbb{R}^{|\mathcal{S}|}}L(V,\rho)\leq\max_{\rho\in H}\min_{V\in\mathbb{R}^{|\mathcal{S}|}}L(V,\rho)=L(V^{\star},\rho^{\star})$. The equality only holds for $\rho=\rho^{\star}$ due to the unique saddle-point as established in Corollary 2. Define $\Gamma(\rho(t)):=\nabla_{\rho}L(\lambda(\rho(t)),\rho(t))$ and with a slight abuse of notation $\zeta(\rho(t)):=\zeta(\rho(t),\Gamma(\rho(t)))$. Then,
@@
Lastly, we address the stability of the iterate sequence, property (A5). The dual variable is constrained to the compact set $H$. We, therefore, only have to ensure that $\sup_{k}\|V_{k}\|<\infty$ almost surely. We apply the scaling limit method outlined in Borkar and Meyn (2000) and Borkar (2023, Theorem 4.1), which states that the global asymptotic stability of the origin for a certain scaling limit ODE is a sufficient condition for the almost sure boundedness of the iterates. The scaling limit for the value iterates for a fixed occupancy $\rho$ is defined as
@@ Appendix B Asynchronous Almost Sure Convergence: Proof Details
This section contains the proofs of the asynchronous algorithm’s properties listed in Proposition 12. The proofs are presented in a series of lemmas. For (B4) the proof of the synchronous convergence analysis still applies. The two options regarding the behavioral policy update only need to be addressed separately for property (B5).
Lemma 21
@@
Proof  We first derive the conditional expectation of the buffer-based transition estimator. Fix $(s,a)\in\mathcal{D}_{\mathrm{inc}}(s_{k})$, such that at iteration $k$ there is a draw from the corresponding buffer (otherwise $\widehat{\mathcal{P}}_{k}(\cdot|s,a)\equiv 0$ by convention and the claims are trivial). Recall that $\xi_{k}(s,a)$ is drawn uniformly from the list $\mathcal{D}_{k}(s,a)$ and $\widehat{\mathcal{P}}_{k}(s^{\prime}|s,a)=\mathbf{1}\{\xi_{k}(s,a)=s^{\prime}\}$. If $(s,a)=X_{k}=(s_{k-1},a_{k-1})$, then $\mathcal{D}_{k}(s,a)=\mathcal{D}_{k-1}(s,a)\cup\{s_{k}\}$ and $\nu_{k}(s,a)=\nu_{k-1}(s,a)+1$. Conditionally on $\mathcal{F}_{k}^{-}=\sigma(\mathcal{F}_{k-1},s_{k},a_{k})$, the multiset $\mathcal{D}_{k}(s,a)$ is fixed and the uniform draw satisfies
@@
The bounds on the second moments of the martingale differences, (B1)-b, can be derived similarly to the proof of Proposition 7 upon noting that $\|\widehat{\mathcal{P}}_{k}(\cdot|s,a)\|_{2}\leq 1$ and
@@
Moreover, boundedness of $\rho_{k-1}$ and $\|\mathcal{P}_{\mathcal{D}_{k}}(\cdot|s,a)\|_{2}\leq 1$ imply that $\{\mathcal{E}_{k}\}$ is bounded uniformly. This proves (B1)-c.
$\blacksquare$
Lemma 22
@@
Proof  The properties (B2) and (B3) regarding the equilibria of the dynamical systems, hold for the asynchronous projected stochastic ascent-descent scheme because we can relate the asymptotic behavior of any solution to the non-autonomous rescaled ODEs (18) to the limiting ODEs of the synchronous setting (13) following (Borkar, 2023, Section 6.4).
@@
Let $\rho\in H$ be a fixed arbitrary dual variable. Let $\epsilon>0$ and $\Lambda^{(1)}(\cdot)$ be a measurable function with $\Lambda^{(1)}(t)\in\Omega^{\epsilon}_{|\mathcal{S}|}$ for all $t$. By definition, $\Lambda^{(1)}(t)$ is a diagonal matrix for each $t$ with bounded diagonal entries $w_{i}(t)\in[\epsilon,1]$. The measurability and boundedness of $t\mapsto\Lambda^{(1)}(t)$ together with the Lipschitz continuity of $V\mapsto\nabla_{V}L(V,\rho)$ imply that the vector field $\Lambda^{(1)}(t)\nabla_{V}L(V,\rho)$ is measurable in $t$ and locally Lipschitz in $V$. Hence, for every initial condition, the ODE (18a) admits a (unique) Carathéodory solution; see, e.g., standard results in differential equations (Filippov, 2013).
@@
Moreover, equality holds if and only if $\nabla_{V}L(V(t),\rho)=0$, that is, $V(t)=\lambda(\rho)$. By $\eta_{V}$-strong convexity of $V\mapsto L(V,\rho)$, the Polyak-Łojasiewicz inequality holds: $\|\nabla_{V}L(V,\rho)\|_{2}^{2}\geq 2\eta_{V}\,\Phi_{\rho}(V)$ for all $V$. Therefore,
@@
For (B3) the same arguments hold with the unchanged Lyapunov function used in the proof of condition (A4), see (36), since multiplying the vector field by a diagonal $\Lambda^{(2)}(t)\in\Omega^{\epsilon}_{|\mathcal{X}|}$ preserves negativity of the Lyapunov derivative. $\blacksquare$
Lemma 23
@@
Proof  Let Assumption 9 hold. In Algorithm 2, conditional on the history $\mathcal{F}_{k}$, the next state-action pair $(s_{k+1},a_{k+1})$ is generated by first sampling $s_{k+1}\sim\mathcal{P}(\cdot|s_{k},a_{k})$ and then sampling $a_{k+1}\sim\pi_{k}^{b}(\cdot|s_{k+1})$, where the behavioral policy $\pi_{k}^{b}$ is either fixed or (in the on-policy exploration variant) a deterministic function of $\rho_{k}$ via (3), optionally $\epsilon_{k}$-mixed with the uniform policy. Therefore, for any $(s,a),(s^{\prime},a^{\prime})\in\mathcal{X}$,
@@ Appendix C Asynchronous Convergence Rate: Proof Details
This section contains the proof details of the convergence rate theorem, Theorem 16. The first section addresses Remark 14 and Proposition 15, followed by the proofs of Propositions 18 and 19 and finally some additional technical lemmata collected in the third subsection.
C.1 Proofs Related to Assumption 13 and Proposition 15
@@
We first address the sufficient condition for the uniform geometric ergodicity with common constants of the family of Markov chains induced by the dual iterate.
Lemma 24 (Deterministic mixing policy implies Assumption 13)
Assume that
there exists a deterministic state-feedback policy $\pi_{d}:\mathcal{S}\to\mathcal{A}$ such that the resulting Markov chain on $\mathcal{S}$ with transition kernel
$\mathcal{P}_{d}(s^{\prime}|s):=\mathcal{P}\bigl(s^{\prime}|s,\pi_{d}(s)\bigr)$
is uniformly geometrically ergodic. Then Assumption 13(i) and (ii) hold. Specifically, let $M\in\mathbb{N}$, $\epsilon_{0}\in(0,1]$, and a probability measure $\upsilon$ on $\mathcal{S}$ be such that
$\mathcal{P}_{d}^{\,M}(\cdot|s)\geq\epsilon_{0}\,\upsilon(\cdot),\qquad\forall s\in\mathcal{S}.$
Further define
$\pi_{\min}:=\frac{C^{L}}{|\mathcal{A}|\,C^{U}},\qquad\epsilon:=\pi_{\min}^{M}\epsilon_{0},\qquad\varrho:=(1-\epsilon)^{1/M}.$
Then the following hold.
  1. (a)
    For every $\rho\in H$, the state-action kernel $\mathcal{P}_{\rho}$ has a unique stationary distribution $\mu_{\rho}$, and
    $\sup_{x\in\mathcal{X}}\|\delta_{x}\mathcal{P}_{\rho}^{k}-\mu_{\rho}\|_{\mathrm{TV}}\leq C_{\mathcal{X}}\varrho^{k},\qquad k\geq 0,$
    with the common constant $C_{\mathcal{X}}:=\varrho^{-M}$.
  2. (b)
    For every sequence $\{\rho_{t}\}_{t\in\mathbb{N}}\subset H$,
    $\sup_{k\geq 0}\delta\!\left(\prod_{t=k}^{k+m_{\star}-1}\mathcal{P}_{\rho_{t}}\right)\leq\kappa,\qquad m_{\star}:=M+1,\quad\kappa:=1-\epsilon.$
@@
Proof  By the equivalence of uniform geometric ergodicity and the multi-step Doeblin condition, see Meyn and Tweedie (2012, Theorem 16.0.2), there exist $M\in\mathbb{N}$, $\epsilon_{0}\in(0,1]$, and a probability measure $\upsilon$ on $\mathcal{S}$ such that
$\mathcal{P}_{d}^{\,M}(\cdot|s)\geq\epsilon_{0}\,\upsilon(\cdot),\qquad\forall s\in\mathcal{S}.$
For $\rho\in H$, let $\pi_{\rho}$ denote the dual-induced policy (3) and define the corresponding state-transition kernel
$\tilde{\mathcal{P}}_{\rho}(s^{\prime}|s):=\sum_{a\in\mathcal{A}}\pi_{\rho}(a|s)\mathcal{P}(s^{\prime}|s,a).$
Since $C^{L}\leq\rho(s,a)\leq C^{U}$ for all $(s,a)\in\mathcal{S}\times\mathcal{A}$, we have the uniform lower bound
$\pi_{\rho}(a|s)=\frac{\rho(s,a)}{\sum_{b\in\mathcal{A}}\rho(s,b)}\geq\frac{C^{L}}{|\mathcal{A}|\,C^{U}}=\pi_{\min},\qquad\forall\rho\in H,\ s\in\mathcal{S},\ a\in\mathcal{A}.$
Hence, for every $\rho\in H$ and $s,s^{\prime}\in\mathcal{S}$,
$\tilde{\mathcal{P}}_{\rho}(s^{\prime}|s)$ $=\sum_{a\in\mathcal{A}}\pi_{\rho}(a|s)\mathcal{P}(s^{\prime}|s,a)$
$\geq\pi_{\rho}(\pi_{d}(s)|s)\,\mathcal{P}\bigl(s^{\prime}|s,\pi_{d}(s)\bigr)$
$\geq\pi_{\min}\,\mathcal{P}_{d}(s^{\prime}|s).$
@@
By monotonicity under composition, iterating the above inequality $M$ times yields
$\tilde{\mathcal{P}}_{\rho}^{\,M}(\cdot|s)\geq\pi_{\min}^{M}\,\mathcal{P}_{d}^{\,M}(\cdot|s)\geq\pi_{\min}^{M}\epsilon_{0}\,\upsilon(\cdot)=\epsilon\,\upsilon(\cdot),\qquad\forall\rho\in H,\ s\in\mathcal{S}.$
Therefore, for every $\rho\in H$, the kernel $\tilde{\mathcal{P}}_{\rho}^{\,M}$ satisfies the same Doeblin minorization, and therefore
$\delta(\tilde{\mathcal{P}}_{\rho}^{\,M})\leq 1-\epsilon,\qquad\forall\rho\in H.$
We first use this to obtain uniform geometric ergodicity of the frozen state kernels. Let $k=mM+r$ with $m\in\mathbb{N}_{0}$ and $0\leq r<M$. For any probability measures $\eta,\eta^{\prime}$ on $\mathcal{S}$, by submultiplicativity of Dobrushin’s coefficient and the bound $\delta(Q)\leq 1$ for any Markov kernel $Q$,
$\|\eta\tilde{\mathcal{P}}_{\rho}^{k}-\eta^{\prime}\tilde{\mathcal{P}}_{\rho}^{k}\|_{\mathrm{TV}}\leq(1-\epsilon)^{m}\|\eta-\eta^{\prime}\|_{\mathrm{TV}}\leq\varrho^{-(M-1)}\varrho^{k}\|\eta-\eta^{\prime}\|_{\mathrm{TV}}.$
In particular, each $\tilde{\mathcal{P}}_{\rho}$ has a unique stationary law $\tilde{\mu}_{\rho}$ on $\mathcal{S}$, and
$\sup_{s\in\mathcal{S}}\|\delta_{s}\tilde{\mathcal{P}}_{\rho}^{k}-\tilde{\mu}_{\rho}\|_{\mathrm{TV}}\leq\varrho^{-(M-1)}\varrho^{k},\qquad k\geq 0.$
We now lift this to the state-action chain on $\mathcal{X}=\mathcal{S}\times\mathcal{A}$. Define the kernels
$\mathcal{T}(s^{\prime}|s,a):=\mathcal{P}(s^{\prime}|s,a),\qquad\mathcal{B}_{\rho}\bigl(s,a\mid s^{\prime}\bigr):=\mathbf{1}_{\{s=s^{\prime}\}}\pi_{\rho}(a|s).$
Then
$\mathcal{P}_{\rho}=\mathcal{T}\mathcal{B}_{\rho},\qquad\tilde{\mathcal{P}}_{\rho}=\mathcal{B}_{\rho}\mathcal{T}.$
Therefore, if $\tilde{\mu}_{\rho}$ is stationary for $\tilde{\mathcal{P}}_{\rho}$, then
$\mu_{\rho}:=\tilde{\mu}_{\rho}\mathcal{B}_{\rho}$
is stationary for $\mathcal{P}_{\rho}$, that is,
$\mu_{\rho}(s,a)=\tilde{\mu}_{\rho}(s)\pi_{\rho}(a|s).$
Moreover, for any $x\in\mathcal{X}$ and any $k\geq 1$,
$\delta_{x}\mathcal{P}_{\rho}^{k}=(\delta_{x}\mathcal{T})\tilde{\mathcal{P}}_{\rho}^{\,k-1}\mathcal{B}_{\rho}.$
Using non-expansiveness of total variation under Markov kernels, we obtain
$\|\delta_{x}\mathcal{P}_{\rho}^{k}-\mu_{\rho}\|_{\mathrm{TV}}$ $=\|(\delta_{x}\mathcal{T})\tilde{\mathcal{P}}_{\rho}^{\,k-1}\mathcal{B}_{\rho}-\tilde{\mu}_{\rho}\mathcal{B}_{\rho}\|_{\mathrm{TV}}$
$\leq\|(\delta_{x}\mathcal{T})\tilde{\mathcal{P}}_{\rho}^{\,k-1}-\tilde{\mu}_{\rho}\|_{\mathrm{TV}}$
$\leq\varrho^{-(M-1)}\varrho^{k-1}=\varrho^{-M}\varrho^{k}.$
Since $\|\delta_{x}-\mu_{\rho}\|_{\mathrm{TV}}\leq 1\leq\varrho^{-M}$ for $k=0$, this proves assertion (a) with $C_{\mathcal{X}}:=\varrho^{-M}$.
It remains to prove assetion (b). Fix an arbitrary sequence $\{\rho_{t}\}_{t\in\mathbb{N}}\subset H$ and a block starting at time $i$. By the pointwise bound $\tilde{\mathcal{P}}_{\rho_{t}}\geq\pi_{\min}\mathcal{P}_{d}$ derived above, monotonicity under composition gives